Please show me how to solve. Thanks!
2/x-8 + 4/x+8

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

without an equal sign, there is no solution

Answer by stanbon(75887)

2/(x-8) + 4/(x+8)
lcd = (x-8)(x+8)
---
=2(x+8)/lcd + 4(x-8)/lcd
=[2x+16+4x-32]/lcd
=[6x-16]/lcd
=[2(x-8)]/lcd
=2/(x+8)
=================
Cheers,
Stan H.